Grade 11Mechanics

the motion of a body falling from rest in a resisting medium is described by the equation dv/dt=a-bv where a and b are constants the velocity at any time t is given by

Praveen Reddy
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
Dear Student,
dv/dt=a-bv
=>〖dv=〖(a-bv)dt〗〗
=> v=at-bvt
=>v=at/(1+bt) m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)